OpenStudy (anonymous):

Guys i have another one cx Solve the equation x-2y= -9 2x+7y=4

hero (hero):

Hint: x = 2y - 9 Use substitution

OpenStudy (anonymous):

@Hero here's what i have so far 2x-4y=-18 2x+7y=4 3y=-14 y=-4.6 but im not sure thats right .-.

hero (hero):

How did you get 3y = -14?

OpenStudy (anonymous):

-4+7 = 3 -18+4= -14

OpenStudy (anonymous):

@Hero

hero (hero):

Well, but you have to perform the same operation for each term You subtracted 2x - 2x to get zero You have to also subtract 7y - (-4y) 4 - (-18)

hero (hero):

If you did that you would have 11y = 22

OpenStudy (anonymous):

so it'd be y= -4.6 and x=2 @Hero right???

hero (hero):

I get y = 2

hero (hero):

y = 22/11 = 2

OpenStudy (anonymous):

oh i see.. thank you!

hero (hero):

Sorry y = 2 so 2x+7y=4 2x + 7(2) = 4 2x + 14 = 4 2x = 4 - 14 2x = -10 x = -10/2 x = -5